OpenStudy (anonymous):

(-1, 1) and (-3,-5) A. y + 3 = 3(x + 5) B. y = 3x + 4 C. y - 1 = 3(x + 1) D. y + 5 = 3(x + 3)

OpenStudy (anonymous):

(-1, 1) and (-3,-5) Gradient = (y2-y1) / (x2-x1) = (-5-1) / (-3--1) = (-6) / (-2) = 3 Equation of line: y - y1 = m (x-x1), where x1y1 are any point on the line y - 1 = 3 (x+1)

OpenStudy (anonymous):

thks man what about (-3,5)?

OpenStudy (anonymous):

(-3,-5)